OpenStudy (anonymous):

Solve. –3(5 – 15n) = 15(3n – 2) A. 30 B. –15 C. no solution D. all real numbers

OpenStudy (anonymous):

-15 + 45n = 45 n - 30 45n = 45n -45 0n = -45 so C.